# AVA-Bench: Atomic Visual Ability Benchmark for Vision Foundation Models

### v2

@@ Appendix B Experiment Details
The hyperparameters of Stage-1 vision-language alignment pretraining, Stage-2 visual instruction tuning and Stage-3 AVA-Bench evaluation are show in Table 3.
@@ B.1 Metric Details
We use CIEDE2000 luo2001development to calculate the color differences using the colour Python library. Specifically, we convert the predictions and ground-truths from CIE XYZ tristimulus format to CIE L*a*b* colour space with colour.XYZ_to_Lab, followed by the colour.delta_E(pred, gt, method="CIE 2000") for color differences.
@@
where N is the number of test samples, $y_{i}$ is the ground-truth (depth or count) for sample $i$ and $\hat{y}_{i}$ the model prediction.
@@
We use the Generalized Intersection-over-Union (GIoU rezatofighi2019generalized , Equation 2):
@@
where $A$ and $B$ are the prediction and ground-truth bounding boxes and $C$ is the smallest (axis–aligned) enclosing box of $A\cup B$.
@@
We evaluate OCR performance with Average Normalized Levenshtein Similarity (ANLS) biten2019scene :
$\mathrm{NLS}(p,g)$ $=1-\frac{\mathrm{Lev}(p,g)}{\max\bigl{(}|p|,|g|\bigr{)}},$ (3)
$\mathrm{ANLS}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathrm{NLS}(p_{i},g_{i})=\frac{1}{N}\sum_{i=1}^{N}\Biggl{(}1-\frac{\mathrm{Lev}(p_{i},g_{i})}{\max\bigl{(}|p_{i}|,|g_{i}|\bigr{)}}\Biggr{)}.$
@@
Appendix C More Results and Analyses
@@ C.1 Detailed Overall Results
SigLIP-2 0.0843 97.38 81.18 0.225 0.6738 88.19 90.17 72.60 45.33 99.17 58.39 79.2 94.08 12.25 2.4
3 2 1 1 1 1 3 3 2 4 5 4 2 2
AIMv2 0.1008 95.71 62.44 0.254 0.5896 85.04 91.78 72.86 43.61 98.93 60.68 79.71 94.11 19.61 3.9
9 8 3 3 5 4 1 2 3 5 1 3 1 6
SigLIP-1 0.0953 96.58 80.3 0.229 0.6103 87.84 90.94 73.4 45.76 98.66 59.6 78.91 93.66 12.64 3.6
6 7 2 2 4 2 2 1 1 8 4 5 3 3
CLIP 0.08461 97.04 60.44 0.290 0.5787 85.02 86.83 72.32 41.59 99.28 60.42 77.77 93.25 19.85 4.9
4 4 6 7 7 5 4 5 4 3 2 6 5 7
InternVL-2.5 0.08212 97.00 60.88 0.269 0.5850 83.19 72.83 71.63 36.13 99.31 59.99 75.71 91.21 20.09 5.4
2 5 5 5 6 7 7 6 7 2 3 7 7 8
RADIOv2.1 0.07645 97.92 62.44 0.257 0.6617 84.90 85.44 72.50 38.77 99.48 56.59 83.71 93.32 17.16 3.6
@@ C.2 Detailed Analyses for Each AVA
We split localization testing samples based on normalized bounding box sizes (relative to image size), where 0.1 indicates an object occupies 10% of the image area. As illustrated in Figure 10, VFMs surprisingly exhibit minimal performance differences when localizing large objects (0.3–0.5). Conversely, performance disparities amplify as object size decreases, revealing significant weaknesses in MiDas and SAM for smaller objects. Consequently, the lower overall performance of MiDas and SAM is predominantly due to poor handling of small targets. Practitioners should thus consider object size distributions when selecting VFMs; SAM and MiDas remain viable if target objects are predominantly large .
@@
Fine-grained classification results are consistently robust across datasets, with the exception of SAM and MiDas, both of which are known to lack semantically rich features espinosa2024there ; chen2023segment , resulting in poorer performance (see Figure 28).
@@
Appendix D Related Works
D.1 VFM Evaluation
Existing evaluation VFM protocols generally fall into two categories. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification han2022survey and COCO for detection or segmentation thisanke2023semantic . For example, DINOv2 oquab2023dinov2 is evaluated on image and video classification, instance recognition, image retrieval, semantic segmentation, and depth estimation. For each task, a task-specific head is trained.
To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks liu2023visual ; zhu2023minigpt ; chowdhery2023palm . For example, in addition to the traditional task-specific evaluation, AIMv2 fini2024multimodal and RADIO  ranzinger2024radio follow the LLM-based evaluation and use a Llama-3.0(8B) grattafiori2024llama and a Vicuna-1.5(7B) zheng2023judging respectively, demonstrating a shift towards generalized multimodal evaluation.
D.2 Atomic Visual Abilities
@@
A recent work chae24decomposing introduced AVSBench to evaluate whether MLLMs understand basic geometric features, including angle, boundary, orthogonality, and curvature, which they refer to as atomic visual skills. However, AVSBench primarily targets geometric comprehension abilities required for geometric diagrams arising in high-school level mathematics. Moreover, AVSBench provides only test data for MLLMs without addressing potential mismatches between training and test data distributions—an issue highlighted in Section 2.2. Consequently, mispredictions in AVSBench evaluations may result from data distribution mismatches rather than genuine visual deficiencies in VFM. In contrast, AVA-Bench explicitly emphasizes atomic visual abilities essential for general visual reasoning tasks commonly encountered in real-world scenarios. By aligning training and evaluation data distributions, AVA-Bench ensures that evaluation outcomes reliably reflect genuine visual perceptual capabilities of VFMs.
@@
Additionally, a concurrent work wu2025compact defines a set of atomic visual capabilities analogous to ours. However, their goal fundamentally differs from ours: wu2025compact aims to build a visual compositional tuning data recipe that builds complex capabilities from simple atomic capabilities, which can significantly reduce instruction-tuning data volume while maintaining strong performance. In contrast, AVA-Bench’s objective is to systematically evaluate VFMs against atomic visual abilities, pinpointing their exact strengths and weaknesses, and providing a comprehensive diagnostic tool to advance the continual development mai2022online ; lomonaco2022cvpr ; mai2021supervised ; shim2021online ; mai2020batch of robust vision foundation models.

### v3

@@ Appendix B Experiment Details
B.1 Hyperparameter Details
To ensure reproducibility and fairness, we carefully followed the official TinyLLaVA hyperparameter recommendations for stages 1 and 2, maintaining both the global batch size and learning rate as prescribed (in Table 4.). For stage 3, which incorporates LoRA-based fine-tuning, we selected a learning rate of 1e-4 and explored multiple LoRA dimensions (64, 128, 256). To validate these choices, we conducted preliminary experiments using three representative VFMs: DINOv2, CLIP, and SigLIP-2. We evaluated performance on two representative AVA tasks (OCR and Recognition), as summarized in Table 3.
The results consistently show that the recommended learning rate of 1e-4 yields stable and strong performance, whereas alternative learning rates often underperform or lead to instability. Similarly, LoRA dimensions between 64 and 128 produce comparable and reliable results, while extreme values show diminishing returns. Based on these observations, we adopt the recommended configuration (learning rate 1e-4, LoRA dimension 128) throughout our experiments. The overall hyperparameters of Stage-1 vision-language alignment pretraining, Stage-2 visual instruction tuning and Stage-3 AVA-Bench evaluation are shown in Table 4.
Task Model lr 1e-5 lr 1e-4 lr 5e-4 LoRA 64 LoRA 128 LoRA 256
OCR DINOv2 7.26 9.97 10.6 10.85 9.97 10.98
SigLIP-2 79.68 81.18 77.63 81.51 81.18 81.25
CLIP 54.23 60.44 60.9 60.79 60.44 61.73
Recognition DINOv2 83.92 86.31 unstable 86.39 86.31 86.46
SigLIP-2 87.04 88.19 unstable 88.42 88.19 88.36
CLIP 83.08 85.02 unstable 84.79 85.02 84.18
Table 3: Hyperparameter exploration for OCR and Recognition tasks using three representative VFMs. Results are reported across different learning rates and LoRA dimensions.
@@ B.2 Metric Details
We use CIEDE2000 [58] to calculate the color differences using the colour Python library. Specifically, we convert the predictions and ground-truths from CIE XYZ tristimulus format to CIE L*a*b* colour space with colour.XYZ_to_Lab, followed by the colour.delta_E(pred, gt, method="CIE 2000") for color differences.
@@
where N is the number of test samples, $y_{i}$ is the ground‐truth (depth or count) for sample $i$ and $\hat{y}_{i}$ the model prediction.
@@
We use the Generalized Intersection-over-Union (GIoU [76], Equation 2):
@@
where $A$ and $B$ are the prediction and ground‐truth bounding boxes and $C$ is the smallest (axis–aligned) enclosing box of $A\cup B$.
@@
We evaluate OCR performance with Average Normalized Levenshtein Similarity (ANLS) [4]:
$\mathrm{NLS}(p,g)$ $=1-\frac{\mathrm{Lev}(p,g)}{\max(|p|,|g|)},$ (3)
$\mathrm{ANLS}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathrm{NLS}(p_{i},g_{i})$
$=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{\mathrm{Lev}(p_{i},g_{i})}{\max(|p_{i}|,|g_{i}|)}\right).$
@@
Appendix C More Results and Analysis
Figure 20: Examples of Relative Depth AVA Samples.
@@ C.1 Detailed Overall Results
SigLIP-2 0.0843 97.38 81.18 0.225 0.6738 88.19 90.17 72.60 45.33 99.50 58.39 79.2 94.08 12.25 2.4
3 2 1 1 1 1 3 3 2 2 5 4 2 2
AIMv2 0.1008 95.71 62.44 0.254 0.5896 85.04 91.78 72.86 43.61 99.13 60.68 79.71 94.11 19.61 3.9
9 8 3 3 5 4 1 2 3 6 1 3 1 6
SigLIP-1 0.0953 96.58 80.3 0.229 0.6103 87.84 90.94 73.4 45.76 99.07 59.6 78.91 93.66 12.64 3.6
6 7 2 2 4 2 2 1 1 7 4 5 3 3
CLIP 0.08461 97.04 60.44 0.290 0.5787 85.02 86.83 72.32 41.59 99.25 60.42 77.77 93.25 19.85 4.9
4 4 6 7 7 5 4 5 4 5 2 6 5 7
InternVL-2.5 0.08212 97.00 60.88 0.269 0.5850 83.19 72.83 71.63 36.13 99.5 59.99 75.71 91.21 20.09 5.4
2 5 5 5 6 7 7 6 7 4 3 7 7 8
RADIOv2.1 0.07645 97.92 62.44 0.257 0.6617 84.90 85.44 72.50 38.77 99.69 56.59 83.71 93.32 17.16 3.6
@@ C.2 Detailed Analyses for Each AVA
We split localization testing samples based on normalized bounding box sizes (relative to image size), where 0.1 indicates an object occupies 10% of the image area. As illustrated in Figure 8 (b), VFMs surprisingly exhibit minimal performance differences when localizing large objects (0.3–0.5). Conversely, performance disparities amplify as object size decreases, revealing significant weaknesses in MiDas and SAM for smaller objects. Consequently, the lower overall performance of MiDas and SAM is predominantly due to poor handling of small targets. Practitioners should thus consider object size distributions when selecting VFMs; SAM and MiDas remain viable if target objects are predominantly large.
@@
Fine-grained classification results are consistently robust across datasets, with the exception of SAM and MiDas, both of which are known to lack semantically rich features [19, 9], resulting in poorer performance (see Figure 26).
@@
Appendix D VFM Details
Table 6 provides a detailed overview of the vision foundation models (VFMs) evaluated in our study. For each model, we list its architecture, parameter count, and training data, and we further summarize the training methodology in terms of supervision type, process, and loss functions.
VFM Architecture Parameters Training Data Training Details
SigLIP-2 Dual-tower ViT encoders for image and text embeddings with MAP pooling layers So400m (400M) WebLI dataset •Supervision: Supervised (image-text pairs) •Process: Pretrained on 40B samples, large-batch training (32k) •Loss: Pairwise sigmoid ITC + captioning/grounding
AIMv2 ViT-based vision encoder with an autoregressive multimodal decoder for patch and token reconstruction AIMv2-Huge (600M) DFN, COYO, HQITP •Supervision: Supervised (multimodal autoregressive) •Process: Pretrain (224 px) → finetune (336/448 px), long training •Loss: Joint reconstruction of image patches and text tokens
SigLIP-1 Dual-encoder with independent ViT and text transformer projecting to a shared embedding space So400m (400M) WebLI dataset •Supervision: Supervised (image-text pairs) •Process: Large-scale pretraining, efficient setup with 32k batch •Loss: Pairwise sigmoid contrastive loss
CLIP Dual-tower model using a ViT image encoder and Transformer text encoder for contrastive alignment ViT-L/14 (428M) Internet-collected dataset •Supervision: Supervised (image-text pairs) •Process: Weeks-long pretraining on multi-GPU/TPU with large batches ($\geq$32k) •Loss: Contrastive InfoNCE
InternVL-2.5 Large multimodal architecture combining a high-capacity ViT encoder with an LLM for image-text reasoning InternVL2.5 (304M) FaceCaption, GQA, ChartQA, Many other datasets •Supervision: Supervised (multimodal LLM) •Process: Pretrained and finetuned on diverse datasets •Loss: Autoregressive next-token + alignment losses
RADIO v2.1 ViT backbone with conditional positional encoding and multi-teacher feature distillation layers RADIO-Huge (653M) DataComp1B dataset •Supervision: Supervised (teacher-student distillation) •Process: 600k steps, AdamW (WD=1e-4), batch scaling law (eff. BS 1024) •Loss: Multi-teacher distillation from CLIP, DINOv2, SAM-H
DINOv2 ViT backbone with patch embedding and projection heads for self-supervised feature representation learning ViT-Large (300M) LVD-142M dataset •Supervision: Self-supervised (no labels) •Process: Teacher-student distillation pipeline with deduplication and retrieval •Loss: Self-distillation contrastive objective
SAM MAE-pretrained ViT-H image encoder paired with a prompt encoder handling points, boxes, masks, and text queries ViT-H (637M) SA-1B dataset •Supervision: Supervised (segmentation masks) •Process: Pretrained encoder with MAE, promptable training on SA-1B •Loss: Segmentation mask prediction loss
MiDaS-3.0 Multi-scale ResNet-based encoder-decoder network designed for monocular depth prediction from single images ResNet-Encoder (123M) DIML Indoor, MegaDepth, ReDWeb, WSVD •Supervision: Supervised (depth ground truth) •Process: Multi-dataset pretraining, 60 epochs, Adam optimizer with different LR for new vs pretrained layers •Loss: Trimmed MAE (20%) + gradient regularizer
Table 6: Details of Vision foundation model (VFM) used, including architecture, parameter scale, training data, and training procedures.
Appendix E Related Works
E.1 VFM Evaluation
Existing evaluation VFM protocols generally fall into two categories. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification [31] and COCO for detection or segmentation [85]. For example, DINOv2 [70] is evaluated on image and video classification, instance recognition, image retrieval, semantic segmentation, and depth estimation. For each task, a task-specific head is trained.
To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks [55, 117, 13]. For example, in addition to the traditional task-specific evaluation, AIMv2 [21] and RADIO  [75] follow the LLM-based evaluation and use a Llama-3.0(8B) [28] and a Vicuna-1.5(7B) [115] respectively, demonstrating a shift towards generalized multimodal evaluation.
E.2 Atomic Visual Abilities
@@
A recent work [7] introduced AVSBench to evaluate whether MLLMs understand basic geometric features, including angle, boundary, orthogonality, and curvature, which they refer to as atomic visual skills. However, AVSBench primarily targets geometric comprehension abilities required for geometric diagrams arising in high-school level mathematics. Moreover, AVSBench provides only test data for MLLMs without addressing potential mismatches between training and test data distributions—an issue highlighted in Section 2.2. Consequently, mispredictions in AVSBench evaluations may result from data distribution mismatches rather than genuine visual deficiencies in VFM. In contrast, AVA-Bench explicitly emphasizes atomic visual abilities essential for general visual reasoning tasks commonly encountered in real-world scenarios. By aligning training and evaluation data distributions, AVA-Bench ensures that evaluation outcomes reliably reflect genuine visual perceptual capabilities of VFMs.
@@
Additionally, a concurrent work [98] defines a set of atomic visual capabilities analogous to ours. However, their goal fundamentally differs from ours: [98] aims to build a visual compositional tuning data recipe that builds complex capabilities from simple atomic capabilities, which can significantly reduce instruction-tuning data volume while maintaining strong performance. In contrast, AVA-Bench’s objective is to systematically evaluate VFMs against atomic visual abilities, pinpointing their exact strengths and weaknesses, and providing a comprehensive diagnostic tool to advance the continual development [61, 57, 62, 79, 60] of robust vision foundation models.
Appendix F Evaluation efficiency
An important advantage of our framework is its efficiency compared to prior LLM-based evaluation protocols. As summarized in Table 7, existing methods typically rely on large language models such as Vicuna-7B and require $\approx$230 A100 GPU hours per vision foundation model (VFM). By contrast, our approach adopts a lightweight 0.5B LLM and smaller training data (1.2M samples in total for stages 1 and 2), which reduces the cost to $\approx$28 A100 GPU hours while still preserving consistent and reliable VFM rankings. This design choice enables practical scaling to a wide range of models without incurring prohibitive resource demands.
Protocol LLM size Stage 1&2 data Stage 1&2 cost Stage 3
Baseline [1] Vicuna-7B 1.9M $\approx$230 A100 h n/a
AVA-Bench Qwen2-0.5B++LoRa(stage 3) 1.2M $\approx$28 A100 h Each AVA: avg 4 A100 h
Table 7: Evaluation Efficiency Table
Dataset Copyright License
Object365 Objects365 Consortium CC By 4.0
LVIS LVIS Consortium CC By 4.0
iNaturalist-2021 iNaturalist (Terms of Service) MIT
DIOR N/A N/A
VQAv2 VQA Consortium CC BY 4.0
FSC CVLab at StonyBrook MIT
CARPK Original image owners (PUCPR/PKLot) N/A
Crowd Surveillance Dataset N/A N/A
CUB-200-2011 Annotations: Catherine Wah et al.; images: original owners CC0 (Public Domain)
FGVC-Aircraft Annotations: S. Maji et al.; images: original owners Research only (Non-commercial)
KITTI Andreas Geiger, Philip Lenz, Christoph Stiller, Raquel Urtasun CC BY-NC-SA 3.0
NYU-DepthV2 N/A N/A
coco-text SE(3) Computer Vision Group, Cornell Tech CC BY 4.0
IIIT5K IIIT Hyderabad (annotations); images: original owners N/A
TextVQA VQA Consortium CC BY 4.0
EgoOrientBench N/A N/A
CURE-OR OLIVES at Georgia Institute of Technology MIT
Moment_int_time Moments in Time authors Research/Educational only
DTD N/A N/A
KTH-TIPS N/A N/A
KTH-TIPS2 N/A N/A
Places365 MIT CSAIL, Bolei Zhou; images: original owners MIT (code); images original copyright owners
AID AID authors (Gui-Song Xia et al.); images from Google Earth providers N/A
RAF-DB N/A N/A
ExpW N/A N/A
Table 8: Dataset copyright and licensing information for all datasets used in AVA-Bench
For stage 3, our framework further leverages LoRA-based fine-tuning, where each AVA is trained on only 6K–10K samples. This procedure requires $\approx$4 A100 GPU hours per AVA on average, making it highly lightweight compared to full model finetuning. In summary, AVA-Bench achieves a more diagnostic evaluation with considerably lower overhead than prior work.
Appendix G Dataset Copyright/License
To ensure ethical and legal use of datasets, we summarize the copyright and licensing information of all benchmarks employed in our experiments (Table 8). The majority of the datasets we use are publicly available under open licenses such as CC BY 4.0, MIT, or CC0, which permit research and redistribution with proper attribution. Some datasets (e.g., FGVC-Aircraft, Moments in Time) are restricted to research-only or educational use, and we adhered to these conditions. For datasets without explicit licensing details, we used them strictly within the scope of non-commercial academic research.